OpenStudy (anonymous):

For the love of God...could someone PLEASE help me with Word Problems - Jerry lives 780 miles from Mary. Jerry leaves his house, bound for Mary's @ the same time Mary leaves her home, bound for Jerry. She is traveling 10-mph faster then Jerry. If they meet after 6 hours, what is the rate they traveled?

OpenStudy (anonymous):

d=rt

OpenStudy (anonymous):

let x = rate at which jerry travels, x+10 = rate at which mary travels. they traveled for 6 hours the total distance of 780 miles so: 6(x) + 6(x+10) = 780 6x + 6x + 60 = 780 12x + 60 = 780 12x = 720 x = 60 So jerry traveled at x = 60 mph, mary = x+10 = 70mph

OpenStudy (anonymous):

THANK YOU THANK YOU TANK YOU - I have been working on this bloody problem for over 2.5 hours now... did I mention THANK YOU!

OpenStudy (anonymous):

np